Two pans of a balance are 67 cm apart. The fulcrum of the balance has been shifted 1.55 cm away from the center by a dishonest shopkeeper. By what percentage is the true weight of the goods being marked up by the shopkeeper? Assume the balance has negligible mass. Answer in units of %.
Answer #1

Consider the correct balance

Take moment about the center

=====================

Consider the faulty balance

Take moment about the center

Mark up = 1.097 -1 =0.097 *100=9.7

ANSWER: =

==============